the moment of inertia of a circular disc about an axis passing through it center and perpendicular to its plane is 10kg m^2it is rotating clockwise with an angular speed 60rpm.another didc with a rotational inertia of 5kgm^2 is rotating anticlockwise with an angular speed 30rpm.at some instant this disc is dropped on top of first disc.calculate final angular speed of the system.
1
Expert's answer
2020-04-08T10:58:48-0400

As per the given question,

Moment of inertia of the first disc "(I_1)=10kg m^2"

Angular speed of the first disc"(\omega_1)=60 rpm"

Moment of inertia of the second disc "(I_2)=5kgm^2"

Angular speed of the second disc "(\omega_2)=30rpm"

When the smaller disc will drop on the larger, then net momentum always be conserve,

Let the final angular momentum be "\omega_3"

Hence, applying the conservation of the angular momentum

"I_1\omega_2-I_2\omega_2= (I_1+I_2) \omega_3"


"\Rightarrow \omega_3= \dfrac{I_1\omega_2-I_2\omega_2}{I_1+I_2}"

"\Rightarrow \omega_3= \dfrac{10\times 60-5\times 30}{10+5}"


"\Rightarrow \omega_3=\dfrac{600-300}{15}"

"\Rightarrow \omega_3=\dfrac{300}{15}=20rpm"